OpenStudy (anonymous):

Factor: x2 + 11x + 10 Answer (x + 10)(x +1) (x + 11)(x + 10) (x + 2)(x + 5) prime

OpenStudy (john_es):

Use the quadratic formula, \[ax^2+bx+c=0\Rightarrow x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (yttrium):

It's factorable. Don't need to use the quadratic formula.

OpenStudy (john_es):

@Ytrium is true. Also you can use the Newton formulas to find the solutions, \[x^2-(x_1+x_2)x+x_1x_2=0\]

OpenStudy (john_es):

As you'll see \[x_1=-10;\ \ x_2=-1\]

OpenStudy (anonymous):